Find a 99% confidence interval for the mean � of a normal population with variance 9

from the sample 30,42,40,34,48,50.


1
Expert's answer
2021-11-24T03:44:58-0500
"mean=\bar{x}=\dfrac{1}{n}\sum _ix_i""=\dfrac{1}{6}(30+42+40+34+48+50)=\dfrac{244}{6}=\dfrac{122}{3}"




"\approx40.6667"

The critical value for "\alpha = 0.01"  is "z_c = z_{1-\alpha/2} = 2.5758."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(\dfrac{122}{3}-2.5758\times\dfrac{3}{\sqrt{6}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(37.512, 43.821)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "37.512 < \mu < 43.821," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(37.512, 43.821)."